OpenStudy (anonymous):

I need help on the following: Solve the following system of equations 7x+3y= -16 3x+5y= 8 x= ? y= ?

OpenStudy (anonymous):

rewrite the equations as 3(7x+3y=16) and 7(3x+5y=8) then subtract to solve for y. Then substitute y into any equation to find x.

OpenStudy (anonymous):

The equation on the top should be -16 21x +9y = -48 21x +35y=56 Bottom minus top 0x +26y = 104 y = 104/26 = 4 21x +9(4) = -48 21x +36 = -48 21x = -84 x = -84/21 = -4 x=-4 y= 4

OpenStudy (anonymous):

Thank you so much :D